40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jun 12 19:41:04 2010, Alberto, Update, Locking, 40m Upgrade Optickle Model optickleIFOworkingpoint.pdf
    Reply  Sun Jun 13 08:28:44 2010, valera, Update, Locking, 40m Upgrade Optickle Model 
       Reply  Mon Jun 14 07:57:07 2010, alberto, Update, Locking, 40m Upgrade Optickle Model 
          Reply  Tue Jun 15 21:28:44 2010, alberto, Update, Locking, 40m Upgrade Optickle Model allTransferFunctions_DARMoff_0.pdfallTransferFunctions_DARM2AS_10pmDARMoffset.pdfJan_DARM2AS.pdf
Message ID: 3072     Entry time: Sat Jun 12 19:41:04 2010     Reply to this: 3074
Author: Alberto 
Type: Update 
Category: Locking 
Subject: 40m Upgrade Optickle Model 

I wrote down the settings according to which I tuned the optickle model of the 40m Upgrade.

Basically I set it so that:

  1. PRC alone anti-resonant for the carrier and resonant for both sidebands
  2. SRC alone resonant for the carrier and resonant for the f2 sideband

In this way when the carrier becomes resonant in the arms we have:

  1. carrier resonant in PRC and anti-resonant in SRC
  2. f1 resonant in PRC and non resonant in SRC
  3. f2 resonant in SRC

The DARM offset for DC readout is optional, and doesn't change those conditions.

I also plotted the carrier and the sideband's circulating power for both recycling cavities.

I'm attaching a file containing more detailed explanations of what I said above. It also contains the plots of field powers, and transfer functions from DARM to the dark port. I think they don't look quite right. There seems to be something wrong.

Valera thought of fixing the problem, removing the 180 degree offset on the SRM, which is what makes the sideband rather than the carrier resonant in SRC. In his model the carrier becomes resonant and the sideband anti-resonant. I don't think that is correct.

The resonant-carrier case is also included in the attachment (the plots with SRMoff=0 deg). In the plots the DARM offset is always zero.

I'm not sure why the settings are not producing the expected transfer functions.

Attachment 1: optickleIFOworkingpoint.pdf  101 kB  | Hide | Hide all
optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf
ELOG V3.1.3-